OpenStudy (anonymous):

Alfonso is 8 years older than Ashley. Five years ago Alfonso was three times as old as Ashley. How old is Ashley now?

OpenStudy (anonymous):

Al = As + 8 (Al-5) = 3(As-5) -------------------- Substitute for Al and solve for Ashley's age. --- ((As+8)-5) = 3As - 15 As+3 = 3As - 15 2As = 18 Ashley = 9 yrs old

OpenStudy (anonymous):

Let's call Alfonso's age x, and Ashley's age y. Right now, x=y+8, from the first given. The second given provides us with x-5=3(y-5). Simplify that to x=3y-10, and subtitute, y+8=3y-10, 2y=18, y=9.

OpenStudy (anonymous):

letx = Ashley letx + 8 = Alofoso x + x + 8 = 3x + 5 so Ashley was 9 yrs old.

OpenStudy (anonymous):

thankyou